Find the χ2 (chi-square) value and P-value from these results. Cite my P-value for your answer.

252 24
208 23

my P-value = ________

(1) My P-value is about 0.238

(2) My P-value is about 0.762

(3) My P-value is about 0.374

(4) My P-value is about 0.626
Answer #1

Given :

Contegency table :

Total
252 24 276
208 23 231
Total 460 47 507

Expected frequency = (row sum * column sum)/Grand sum

E1 = 276*460/507 = 250.14

E2 = 276*47/507 = 25.586

E3 = 231*460/507 = 209.586

E4 = 231*47/507 = 21.414

Chi square goodness of fit test :

Category Observed frequency (Oi) Expected frequency (Ei) (Oi-Ei)^2/Ei
1 252 250.414 0.01
2 24 25.586 0.0983
3 208 209.586 0.012
4 23 21.414 0.1175
Total 507 507 0.238

Test statistic :

\chi^2 = \sum (Oi-Ei)^2/Ei = 0.238

Degree of freedom = (r-1)(c-1) = (2-1)(2-1) = 1*1 = 1

P-value : P-value corresponding to \chi^2 = 0.238 and df = 1 is 0.626

Therefore the χ2 (chi-square) value = 0.238 and P-value = 0.626

Answer - option 4